### v1

# Forgetting Order of Continual Learning: Examples That are Learned First are Forgotten Last

Abstract
Catastrophic forgetting poses a significant challenge in continual learning, where models often forget previous tasks when trained on new data. Our empirical analysis reveals a strong correlation between catastrophic forgetting and the learning speed of examples: examples learned early are rarely forgotten, while those learned later are more susceptible to forgetting. We demonstrate that replay-based continual learning methods can leverage this phenomenon by focusing on mid-learned examples for rehearsal. We introduce Goldilocks, a novel replay buffer sampling method that filters out examples learned too quickly or too slowly, keeping those learned at an intermediate speed. Goldilocks improves existing continual learning algorithms, leading to state-of-the-art performance across several image classification tasks.
1 Introduction
The surge in deploying deep models in real-world scenarios has heightened the need for models capable of adjusting to changing environments. Continual learning (see extensive reviews, e.g [10, 19, 37]), which involves training deep learning models on new data even after mastering previous tasks, is crucial in this context. However, continual learning encounters a significant obstacle called catastrophic forgetting [12, 24, 32]: as models learn from new data, they tend to forget previous tasks, resulting in decreased performance on those tasks. Despite extensive research efforts, catastrophic forgetting persists as a central challenge in continual learning.
We begin by empirically examining catastrophic forgetting, identifying which samples deep models tend to forget when trained on new tasks. Our findings reveal a last-in-first-out pattern, where examples learned later are forgotten more readily than those learned earlier. Fig. 1 visualizes this by showing which examples are likely to be forgotten or remembered over time. This phenomenon is connected to the simplicity bias [48, 50] of neural networks, where networks learn simple examples faster than complex ones. However, complex examples are more likely to be forgotten, while simple examples are more likely to be remembered.
This observation has an important consequence: by measuring how fast examples are being learned by the model, we can predict in advance which examples are more susceptible to forgetting. This holds true regardless of the distribution of the new task introduced to the model. Furthermore, we demonstrate that models showing varying degrees of forgetting do so consistently: models with less catastrophic forgetting recall most of the examples retained by those with more forgetting, with the additional ability to remember slightly more complex examples.
Replay-based continual learning methods use a memory buffer to address catastrophic forgetting, storing a small subset of original task examples. Typically, this buffer is uniformly sampled from previous tasks, and different continual learning techniques vary in how they utilize it. Our findings indicate that while a replay buffer helps mitigate forgetting, a similar pattern emerges: examples learned later are forgotten sooner. Increasing buffer size or improving the underlying model’s architecture can alleviate more of the forgetting to some extent, but the fundamental phenomenon persists. This observation implies that when strengthening a continually trained model, we can predict in advance which examples it will better remember compared to its weaker version.
This raises the question: can one exploit this phenomenon and reduce forgetting by cleverly selecting samples for the replay buffer? Our observations reveal that including in the replay buffer more examples that the model can potentially remember given more practice can improve a large range of continual learning methods. Given the fixed buffer size, these examples replace the fastest-to-learn examples, which the model is likely to remember anyway, and the slowest-to-learn examples, which it is likely to forget. Leveraging these insights, we introduce a practical sampling method named Goldilocks. In constructing the replay buffer, Goldilocks eliminates examples learned too quickly or too slowly, ensuring the buffer comprises examples learned at just the right speed.
Integrating Goldilocks sampling into different competitive continual learning algorithms is computationally negligible, and results in a significant improvement across a diverse range of image classification tasks, surpassing state-of-the-art performance in many cases. Unlike prior replay-buffer sampling methods, which often exhibit effectiveness only within specific contexts or with particular algorithms, Goldilocks consistently performs well across all evaluated settings.
Figure 1: Test examples that networks tend to either remember or forget. Networks were trained on 2 tasks from CIFAR-100. Examples from the first task that were learned quickest (left) and slowest (right) are shown. Slowest examples were often forgotten, while the quickest were seldom forgotten.
1.1 Our contribution
  1. 1.
    We demonstrate consistent catastrophic forgetting across diverse continual learning approaches, highlighting certain examples’ susceptibility to forgetting regardless of the specific continual learning framework.
  2. 2.
    Networks are more likely to forget examples learned during later stages of the learning than examples learned early on. This observation allows us to predict during training the likelihood a given example will be forgotten.
  3. 3.
    We introduce Goldilocks, a novel and efficient replay buffer sampling method. Goldilocks selectively samples examples learned midway through training, discarding examples that are too simple or too complex for the model. Goldilocks reduces catastrophic forgetting and improves the performance of a large variety of continual learning methods.
1.2 Related work
Existing studies primarily focus on mitigating catastrophic forgetting [25, 28, 29, 45, 47], rather than its characterization. Some works have explored catastrophic forgetting from a model perspective: [36, 42] showed differences in forgetting patterns across model layers, while [34, 40] demonstrated the impact of training hyper-parameters. [35] identified task complexity as a factor influencing forgetting rates. Our study differs by examining the relationship between learning speed and forgettability of examples, emphasizing a data-centric perspective.
Simplicity bias. Empirical evidence supports the hypothesis that neural networks initially learn simple models, gaining complexity over time [7, 14, 15, 20, 21, 22, 23, 38, 49, 54]. This phenomenon, known as simplicity bias [11, 48], is extended in our work to catastrophic forgetting, illustrating a reverse simplicity bias where complex examples are forgotten before simpler ones.
Forgetting dynamics. Several studies have directly explored the dynamics of forgetting. [31, 53] revealed that, even in non-continual settings, certain examples tend to be less forgettable during training. [33] demonstrated that the forgetting dynamics of noisy examples differ from those of clean labels and proposed adjusting the rehearsal process accordingly. Our results align with these findings, further delineating which examples are more prone to forgetting and retention, establishing connections between catastrophic forgetting and simplicity bias, and proposing a method to identify examples likely to be forgotten proactively. Additionally, our work illustrates how to leverage these insights to enhance the sampling of the replay buffer for various continual learning approaches.
Replay buffer sampling functions. Replay-based methods mitigate catastrophic forgetting by storing examples from previous tasks in a memory buffer of fixed size. Despite the availability of various sampling strategies [1, 3, 6, 44, 52, 55], their effectiveness often depends on specific conditions or algorithms. As a result, many replay-based methods resort to uniform sampling [5, 16, 25, 30, 41, 43, 46] in practice. Unlike previous methods, Goldilocks can improve replay-based methods across diverse tasks and scenarios.
2 Empirical analysis of catastrophic forgetting
In this section, we empirically study catastrophic forgetting from a data-centric perspective, drawing a connection between the speed at which a model learns examples, and its likelihood to forget them.
2.1 Definitions
We examine the continual learning classification setting, where the training dataset $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ consists of examples $x\in\mathcal{X}$ and their corresponding labels $y\in\mathcal{Y}$. This dataset is divided into $T>1$ tasks, with each task $t\in[T]$ represented as $\mathcal{D}_{t}=\{\mathcal{X}_{t},\mathcal{Y}_{t}\}$, and $\mathcal{D}=\bigcup_{t=1}^{T}{D_{t}}$. There is no overlap between the data of different tasks. A deep model $f:\mathcal{X}\rightarrow\mathcal{Y}$ is sequentially trained on each task $\mathcal{D}_{t}$ from $t=1$ to $t=T$, with $E\in\mathbb{N}$ epochs per task. Each task also has a separate test set from the same distribution. During training on a task $\mathcal{D}_{t}$, the model $f$ is provided with a replay buffer $B$ of fixed-memory size, containing examples from previous tasks ${1,...,t-1}$. Notably, $|B|\ll|\mathcal{D}|$, meaning only a fraction of the data can be stored in the replay buffer. It’s worth mentioning that the buffer may be empty ($B=\emptyset$), representing training $f$ without a memory buffer.
Learning speed. Neural networks demonstrate a trend in learning examples, where certain examples are consistently learned before others across various deep neural models [2, 9, 17]. This orderliness is often connected to the simplicity bias phenomenon [18], as examples learned earlier tend to be simpler than those learned later. Rather than measuring the example’s complexity, which can be defined in many different ways, we use the speed at which examples are learned by the deep model, which can be measured efficiently and directly. Given a model $f$ and an example $(x,y)$ either in the train or test data, we define the learning speed of the example $(x,y)$ to be:
$learning\_speed(x,y)=\frac{1}{E}\sum_{e=1}^{E}\mathds{1}[f_{t}^{e}(x)=y]$ (1)
Here, $f_{t}^{e}$ denotes the intermediate model $f$ after training for $e\in[E]$ epochs on task $\mathcal{D}_{t}$. This definition is based on the accessibility score introduced in [17], adapted to be computed by a single model rather than an ensemble, enabling efficient computation throughout the training of $f$. Intuitively, this score indicates how rapidly the model learns an example: an example correctly classified from the early stages of learning will sustain $f_{t}^{e}(x)=y$ for more epochs, resulting in a higher learning speed.
Computing the learning speed. To calculate the learning speed of each example, we maintain a boolean matrix, called the epoch-wise classification matrix, $M\in\left\{0,1\right\}^{E\times|\mathcal{D}_{t}|}$ throughout $f$’s training on task $\mathcal{D}_{t}$, indicating whether the model correctly classified each example during learning. Specifically, for example $(x_{i},y_{i})\in\mathcal{D}_{t}$, we have $M_{e,i}=\mathds{1}[f_{t}^{e}(x_{i})=y_{i}]$. An example’s learning speed is then the average of the $M$ across epochs. This matrix can be computed during the forward pass of the network, incurring minimal to no additional computation overhead. Furthermore, memory usage can be further reduced by directly computing the matrix’s mean during training.
2.2 Preliminaries
Class incremental vs. task incremental learning. While similar, task incremental learning distinguishes itself from class incremental learning by providing the task identity for each example during testing, enabling the use of the specific classifier of the classes within the task. As a result, performance in class incremental learning is typically much lower. Due to space constraints, our main paper exclusively presents results for task incremental learning. However, findings for class incremental learning, which align qualitatively with the task incremental case, are available in App. A.
Datasets. We investigated various image continual learning classification tasks using split versions of several image datasets, including CIFAR-10, CIFAR-100 [26], and TinyImagenet [27]. The data is split into $T$ tasks by partitioning the classes into $T$ equal-sized subsets. This partitioning is denoted as dataset-T. For example, splitting CIFAR-10 into 5 classes is denoted as CIFAR-10-5, comprising 5 tasks, each with 2 distinct classes. Unless otherwise specified, classes are divided into tasks according to the original order they appeared in the dataset (often alphabetically).
Architectures and training hyper-parameters. In our experiments without a replay buffer, unless stated otherwise, we trained ResNet-18 for $E=100$ epochs per task. We employed a base learning rate of $0.1$ with a cosine scheduler, SGD optimizer, momentum of $0.9$, and weight decay of $0.0005$. All networks were trained on NVIDIA TITAN X. These hyperparameters were chosen arbitrarily based on their performance during joint dataset training and were consistent across all experiments. We anticipate consistent qualitative results despite potential variations in these hyperparameters. When introducing a replay buffer, we employed an experience replay strategy [46], alternating batches of data from the new task and the replay buffer. When comparing against other replay buffer sampling methods, we maintained the same ResNet-18 settings. When integrating with other continual learning algorithms, we evaluated all methods within the framework of [4, 5], using the best hyperparameters suggested in the original papers, while only altering the replay buffer sampling method.
2.3 Correlating learning speed with catastrophic forgetting
(a) Test accuracy
(b) $M$ train
(c) $M$ test
(d) Speed vs. forgetting
[…]
We begin by exploring catastrophic forgetting in a simplistic case, without using replay buffer ($B=\emptyset$), and with only $T=2$ tasks. In Fig. 2(a), we plot the mean test accuracy of each task when training $10$ networks on CIFAR-100-2. The dashed black line marks the transition from the first to the second task. Consistent with prior research, the networks exhibit significant catastrophic forgetting, evidenced by a sharp decline in the accuracy of the first task during training on the second task. Our objective is to characterize those examples where the network successfully classified after the first task but failed after the second task.
To simultaneously track learning speed and forgetting, we maintain the epoch-wise boolean classification matrix $M$ for each network (formally defined in §2.2). Throughout the training of both tasks, we record, for each epoch, whether the network correctly classified each example from the first task. The initial $E=100$ epochs represent classification during the first task, enabling us to observe the learning speed of the different examples, while the subsequent $E=100$ epochs allow us to monitor the forgetting of the first task during training of the second task.
In Figs. 2(b-c), we examine these epoch-wise classification matrices for both the train and test sets, respectively. To aid visualization, instead of plotting the epoch-wise classification matrix for each network, we aggregate together the networks from all runs using majority voting. The rows are sorted by the learning speed of the data samples. Notably, we observe a correlation between learning speed and catastrophic forgetting: examples learned more quickly during the first task tend to remain correctly classified throughout the second task, whereas slower-learned examples are more prone to immediate misclassification upon task transition, making them susceptible to forgetting. Essentially, slower learning speeds of examples correlate with a higher probability of forgetting by deep models.
[…]
While the results presented here pertain to a simplified case, this phenomenon is robust across diverse datasets, architectures, and hyperparameters. For additional experiments, please refer to App. B.
2.4 How the replay buffer changes the catastrophic forgetting dynamics
(a) Small buffer, $4\%$
(b) Medium buffer, $12\%$
(c) Large buffer, $40\%$
(d) Buffer size vs. speed
Figure 3: Impact of the buffer size on forgetting dynamics. (a-c) Similarly to Fig. 2(c), the first task’s binary matrices $M$, when adding replay buffer of varying buffer sizes ($4\%$, $12\%$, and $40\%$). Rows denote examples, and columns denote epochs, indicating correct classifications by the majority of $10$ networks. Rows are sorted by learning speed. In all cases, examples that were learned faster are more likely to be remembered. With bigger buffers, the networks can remember gradually slower-to-learn examples. (d) The mean learning speed of remembered examples of models with different buffer sizes. Models with bigger replay buffers remember slower-to-learn examples.
To evaluate replay-based methods, we incorporate a replay buffer in our experiments. At the end of the first task, we uniformly sample examples to a fixed-memory replay buffer. During training on the subsequent task, we adopt the experience replay technique [46], alternating batch sampling between the new data and the contents of the replay buffer.
Similarly to §2, we trained $10$ networks on CIFAR-100-2, with a replay buffers of varying sizes: $1000$, $3000$, and $10000$ samples corresponding to $4\%$, $12\%$, and $40\%$ of the data size, respectively. In Figs. 3(a-c), we visualize the aggregated epoch-wise classification matrix of the test data for each case. The rows in each matrix are sorted by the learning speed of the examples. As expected, integrating a replay buffer mitigates some catastrophic forgetting, boosting the model performance on the first task, with larger buffers yielding better results. Nevertheless, similar to the scenario without a replay buffer $B=\emptyset$, a relationship exists between learning speed and catastrophic forgetting: networks tend to remember fast-learned examples while forgetting those learned later even when using a uniformly sampled replay buffer. Notably, models with larger buffer sizes remember most of the examples that models with smaller buffer sizes do, while additionally remembering gradually slower-learned examples. Similar results occur when changing the network architecture (see App.B)
We quantify the relationship between the buffer size and remembered examples’ learning speed. Training $10$ networks on CIFAR-100-2 with various buffer sizes, we plot the mean learning speed of examples remembered by the models for each buffer size. A robust correlation emerges ($r=0.966$, $p\leq 10^{-6}$): models with smaller buffer sizes remember quickly learned examples, whereas larger models with larger buffers enable the model also to remember slower-learned. These results are plotted in Fig. 3(d). Similar results were achieved across diverse datasets, architectures, hyperparameters, and continual learning algorithms, see App. B.
3 Effects of different replay buffer compositions on catastrophic forgetting
Replay-based continual learning methods typically sample the replay buffer uniformly, including both quickly and slowly learned examples. In this section, we analyze various replay buffer compositions and find that medium-speed learned examples are the most beneficial for the replay process. These examples, which the model has only a moderate chance of remembering, prove particularly important. One possible explanation is that increasing the buffer size impacts the catastrophic forgetting of these examples the most, suggesting that focusing on them is advantageous.
3.1 The Goldilocks approach – methodology and definitions
To explore different buffer compositions, we remove a combination of the quickest and slowest learned examples from the data and then sample the buffer uniformly from the remaining examples. This allows us to add a bias to the replay buffer toward examples learned at certain speeds. We propose Goldilocks, a novel sampling method, used both in our empirical analysis and as a practical replay buffer sampling technique. The name Goldilocks reflects the method’s focus on examples learned at an optimal speed, akin to the story where Goldilocks prefers things that are "just right."
Formally, Goldilocks is defined by two hyperparameters, $quick$ and $slow$, which specify the percentage of examples learned too quickly or too slowly to be removed from dataset $\mathcal{D}_{t}$. During training of the model $f$ on $\mathcal{D}_{t}$, we maintain the epoch-wise classification matrix $M$ (see § 2.2). At the end of task $t$, examples in $\mathcal{D}_{t}$ are sorted by learning speed, and the top %$\%quick$ and bottom %$\%slow$ are removed. The replay buffer is then uniformly sampled from the remaining examples. Notably, for tasks beyond the initial two, when reducing the replay buffer size, examples can be uniformly sampled out. The pseudocode for Goldilocks sampling is provided in Algorithm 1.
Algorithm 1 Training continual learning method $CL$ with Goldilocks sampling.
Input: task $\mathcal{D}_{t}$, sample size $|B|$, epochs $E$, amount of quick/slow learned examples to remove $q,s$. Output: buffer of size $|B|$
1:$Classification\_Matrix$leftarrow ${0}^{E\times|\mathcal{D}_{t}|}$
2:for $e=1,...,E$ do
3:     Train the model $f$ one epoch using the selected continual learning method $CL$
4:     for $(x_{i},y_{i})\in\mathcal{D}_{t}$ do $\triangleright$ Can be calculated efficiently in the forward pass
5:         Classification_Matrix$[e,i]\leftarrow\mathds{1}[f(x_{i})=y_{i}]$
6:     end for
7:end for
8:$Learning\_Speed$leftarrow Mean($Classification\_Matrix$, axis=0)
9:$\mathcal{D}_{t}^{{}^{\prime}}\leftarrow$ remove $q\%$ highest and $s\%$ lowest $Learning\_Speed$ examples from $\mathcal{D}_{t}$
10:return $|B|$ examples sampled uniformly from $\mathcal{D}_{t}^{{}^{\prime}}$
3.2 Which examples should appear in the replay buffer
(a) CIFAR-100-2, $4\%$ buffer
(b) CIFAR-100-2, $40\%$ buffer
(c) TinyImagenet-2, $20\%$ buffer
Figure 4: Comparing different replay buffer compositions. Each matrix compares the mean final test accuracy of 10 networks trained with different replay buffer compositions achieved by modifying Goldilocks’ $quick$ and $slow$ hyperparameters. The ($quick$, $slow$) = ($0\%$, $0\%$) case, marked by a black box, corresponds to uniform sampling. (a) CIFAR-100-2 with a $4\%$ buffer. Removing faster-to-learn examples is more beneficial. (b) CIFAR-100-2 with a $40\%$ buffer. Removing slower-to-learn examples is more beneficial. (c) TinyImageNet-2 with a $40\%$ buffer. Across all scenarios, focusing on examples learned in the middle of the learning process is most beneficial. Smaller buffers exhibit greater benefits from prioritizing faster-to-learn examples.
(d) (A)–>(B)
(e) (A)–>(C)
Figure 5: Comparing replay buffer compositions when training the same original task, followed by different subsequent tasks. We consider 3 tasks, A, B, and C, training networks on task A followed by either B or C. All tasks are subsets of CIFAR-100, where Task A is classes 0-24, B is classes 25-49 and C is classes 50-74. (a) training on task A followed by B. (b) training on task A followed by C. The results format is similar to Fig. 3.2. Despite changes in the subsequent tasks, effective and ineffective buffer compositions for the original task A remain consistent.
(a) GEM, CIFAR-100-2, $1\%$ buffer
Figure 6: Comparing different replay buffer compositions for GEM with a buffer of $1\%$. The experimental setup replicates the one done in Fig. 3.2. Focusing on examples learned midway through the learning process is most beneficial. Due to the smaller buffer, removing more easier examples is better.
We investigate the effect of constructing the replay buffer with examples learned at different stages of the learning process. This is achieved by adjusting the $slow$ and $fast$ parameters in Goldilocks (see §3.1). Regardless of these parameters’ values, the replay buffer maintains a constant size but is sampled from different data distributions. We evaluate a wide range of $slow$ and $fast$ values by training 10 networks per parameter set on various datasets and buffer sizes. The mean final test accuracy across all tasks is plotted in a grid-like heatmap, in Fig. 3.2. To aid visualization, standard errors are plotted separately, see App. C. Notably, the point $(slow,fast)=(0,0)$ (marked by a black box) represents training with a uniformly sampled replay buffer, as is done in the original works.
In Figs. 3.2(a-b), we compare replay buffer compositions for networks trained on CIFAR-100-2. We consider a small buffer of size $1000$ ($4\%$ of data size) and a large buffer of size $10000$ ($40\%$ of data size). Blank entries denote infeasible parameters where too much data is removed, rendering a $40\%$ buffer impossible. In both cases, removing both slowly- and quickly-learned examples proves beneficial, with parameters mutually influencing each other. Focusing on examples learned midway through the learning process, where the network can either remember or forget, is most effective in reducing catastrophic forgetting. When the buffer is smaller, optimal results are achieved with a bias toward fast-learned examples, as models struggle to remember simpler examples in such scenarios.
In Fig. 4(c), we replicate this experiment with TinyImagenet-2, with a buffer size of $10000$ ($20\%$ of the data). Similarly to the CIFAR-100-2 case, focusing on examples learned midway through learning yields the best performances. It’s worth noting that across all cases, the range of hyperparameters that enhance performance is broad and continuous, suggesting that improvements are not confined to specific hyperparameter sets; instead, various methods of focusing on midway-learned examples yield positive outcomes. Similar results with class incremental learning can be found in App. A.
3.3 Independency of the optimal buffer composition on subsequent tasks
We empirically examine how subsequent tasks impact the ideal composition of a replay buffer for a given task. We find that regardless of the similarity or dissimilarity between subsequent tasks and the original task, the optimal replay buffer composition remains largely independent and consistent. To show this, we conduct experiments with three distinct tasks denoted A, B, and C. By training the model on task A and subsequently introducing either task B or task C while keeping the original task unchanged, we analyze the ideal replay buffer composition under varying subsequent tasks.
We initially focus on tasks from the same datasets. We pick task A to be the first 25 classes of CIFAR-100, task B to be classes 26 to 50, and task C to be classes 51 to 75. In Fig. 5, we depict the mean final accuracy of 10 networks across various replay buffer compositions when using either task B or task C as the subsequent task after task A. Despite some quantitative variations in performance, the qualitative results consistently indicate that good buffer compositions for one case remain effective for the other, and vice versa. These results suggest that the ideal replay buffer composition of task A is robust to differences between tasks B and C in this context.
Further, we explore scenarios involving either changing the classification task or replacing it with memorization by training on random labels. We adopt the RotNet approach [13] to change the classification task. We pick tasks A and B to be different subsets of CIFAR-100, while task C involves a rotation classification task, where each example in task A is randomly rotated between $90\degree,180\degree$, or $270\degree$ degrees, with the task being to classify these rotations regardless of the original label. For the random case, tasks A and B remain unchanged, while task C involves the same subset of task B, but with labels chosen uniformly at random, as suggested in [57]. In both scenarios, we observe consistent results: the ideal replay buffer composition of task A remains robust to differences between tasks B and C in both contexts. Further details and heatmaps for the results are available in App. D.
Choosing hyperparameters. Implementing Goldilocks practically requires identifying the optimal $quick$ and $slow$ hyperparameters without extra data. While simple heuristics can often suffice due to the broad range of hyperparameters yielding good results, our findings propose an alternative approach for hyperparameter fine-tuning. By conducting a grid search for both parameters, training the network on the original task, and using the rotated version as the subsequent task, efficient parameter selection can be achieved. The smooth impact of parameters on the results implies that a coarse grid search is often adequate, reducing computational burden.
3.4 Goldilocks sampling improves various continual learning methods
Many replay-based continual learning algorithms sample the replay buffer uniformly. In this section, we investigate the impact of replacing uniform sampling with Goldilocks in various such algorithms, including DER [5], DER++ [5], X-DER [4], AGEM [8], ER [46], GEM [30], RPC [39], and FDR [51]. Our results demonstrate that Goldilocks consistently improves the performance of all tested continual algorithms, thereby advancing the state of the art in multiple image classification tasks.
We replicate the experimental setup outlined in §3.2, exploring various compositions of the replay buffer in a grid-like fashion. In Fig. 6 we plot a heatmap example for GEM trained on CIFAR-100-2, employing a buffer size of 500 ($2\%$ of the dataset size). We focus on small buffers, as the original works often mainly report results for such buffers. Our findings mirror our previous findings on the experience replay case: a large set of hyperparameters improves the algorithm’s performance, while prioritizing the removal of slowly learned examples proves advantageous given the buffer’s small size. Comparable outcomes across different algorithms, datasets, and class incremental learning scenarios are detailed in Apps. A,E.
Next, we compare continual learning methods with Goldilocks on popular continual learning benchmarks, including CIFAR-10-5, CIFAR-100-5, and TinyImageNet-10 datasets. The $quick$ and $slow$ hyperparameters are determined through RotNet training, as outlined in §3.3. Although these parameters might vary per task when $T>2$, we maintain consistent values across tasks to mitigate an excessive number of hyperparameters. Goldilocks consistently improves various continual learning algorithms, indicating its compatibility with diverse methods. These results are depicted in Fig. 7.
3.5 Comparing Goldilocks to other replay buffer sampling methods
While numerous replay buffer sampling methods have been proposed over the years, uniform sampling remains prevalent in practical implementation within continual learning algorithms. This preference may stem from the specific settings or algorithmic modifications required for other sampling methods to work well. These methods do not consistently improve different continual learning algorithms across datasets and buffer sizes. In contrast, Goldilocks sampling improves a wide range of continual learning algorithms across various buffer sizes and datasets.
We explored various replay buffer sampling strategies, including uniform sampling, max entropy, IPM [56], GSS [1], and Herding [44]. We trained $10$ networks on CIFAR-10-2 and CIFAR-100-2 datasets with buffer sizes of $1000$ and $10000$ (equivalent to $4\%$ and $40\%$ of the dataset size, respectively) using experience replay. Table 3.5 presents the mean test accuracy across all tasks for each case.
Goldilocks consistently outperforms all other sampling strategies across all scenarios, surpassing them in all four cases. In contrast, other sampling methods exhibit efficacy only in specific scenarios. For instance, Herding, which prioritizes examples with mean features close to class mean features, performs better than uniform sampling with smaller buffer sizes but worse with larger buffers. This might be because this preference for typical examples, which are learned early on, is more beneficial with smaller buffers. Conversely, GSS, which selects examples based on gradient distances, outperforms uniform sampling in scenarios with large buffers but performs suboptimally with smaller budgets. This bias towards selecting examples with large gradients favors examples learned slowly, which is more beneficial with large buffers.
(a) CIFAR-100-5
(b) TinyImagenet-10
Figure 7: Goldilocks vs. uniform sampling with various continual learning methods. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error. (a) CIFAR-100-5 with a buffer size of $5\%$. (b) TinyImagenet-10 with a buffer of size $10\%$. Goldilocks significantly improves performance across all methods and datasets.
Table 1: Comparing replay buffer sampling methods. Each entry is the mean test accuracy across all tasks of 10 networks trained using experience replay on CIFAR-100-2. Bold denotes the method with the highest accuracy.
CIFAR-10-2 CIFAR-100-2
Buffer size $4\%$ $40\%$ $4\%$ $40\%$
Random 87.67 95.03 69.03 79.24
Max entropy 84.44 94.63 64.27 77.91
IPM 85.9 94.76 64.16 75.73
GSS 85.28 95.55 64.15 80.11
Herding 88.41 94.53 71.07 78.97
Goldilocks 88.99 96.03 71.53 80.62
4 Conclusion
Limitations. Goldilocks necessitates maintaining an epoch-wise classification matrix to compute the learning speed of each example. This estimation’s accuracy increases with the number of epochs completed. Consequently, in stream-based approaches, where each data point is encountered only once, Goldilocks cannot be applied. In such cases, alternative methods for approximating example complexity may be viable but fall beyond the scope of this study, leaving them for future work.
Another limitation arises in the potential lengthy computation time when fine-tuning hyperparameters in multi-task settings, as the $quick$ and $slow$ hyperparameters may differ between each subsequent task. However, in practice, these hyperparameters can be selected using simple heuristics and kept constant across tasks, as demonstrated in our training on multiple tasks (see Fig. 7).
Discussion. We show a correlation between the speed at which examples are learned and the likelihood of models forgetting these examples when learning new data. This correlation aligns closely with the simplicity bias, suggesting that models tend to forget more complex examples. Although measuring learning speed is a simple proxy for example complexity, its simplicity underscores the potential for developing more sophisticated methods to mitigate catastrophic forgetting, offering avenues for future research.
While most replay-based continual learning works focus on better utilization of the replay buffer, our study demonstrates the potential for buffer improvement, extending the scope of continual learning research beyond buffer utilization alone. By showing that Goldilocks can be implemented with minimal computational overhead, we highlight its applicability to future advancements in continual learning. Furthermore, this work underscores the mutual benefits of concurrent progress in better sampling of the replay buffer and improvements in buffer utilization, suggesting promising directions for future research.
Appendix
Appendix A Class incremental learning
In continual learning, a distinction is made between class incremental learning and task incremental learning. The key difference lies in test time: in task incremental learning, the model knows the task identity of the given examples, allowing it to use a specialized classifier that differentiates between classes within that task. Conversely, in class incremental learning, the model lacks task identity information at test time and must identify both the task and the class of the example, typically resulting in lower performance. Due to space constraints, we report only the task-incremental results in the main paper, as it is the more common benchmark. However, we also tested for class incremental settings in all our experiments. While the performance is lower, the qualitative outcomes remain consistent: buffer compositions that are good for task incremental learning are also good for class incremental settings and vice versa, and Goldilocks significantly enhances a variety of continual learning algorithms in both scenarios.
In Fig.8(a), we extend the results from Fig.4(b) to class incremental settings. We examine different buffer compositions by adjusting the $quick$ and $slow$ hyperparameters of Goldilocks and plot the mean final test accuracy across all tasks for 10 networks trained on CIFAR-100-2 with a $40\%$ buffer, using experience replay. The qualitative results are consistent: the same buffer compositions are effective for both learning scenarios. In Fig.8(b), we replicate Fig.6, applying the same experiment with GEM on CIFAR-100-2 using a $1\%$ buffer, confirming that the results align with those reported in the main paper.
Finally, in Fig.9, we reproduce the results from Fig.7, showing the performance of competitive continual learning algorithms when changing only the buffer sampling strategy to Goldilocks, across CIFAR-100-5, CIFAR-10-5, and TinyImageNet-10. In all datasets and with all continual learning algorithms, Goldilocks significantly improves the performance of the original methods by a significant margin.
(a) CIFAR-100-2, $40\%$ buffer
(b) GEM, CIFAR-100-2, $1\%$ buffer
Figure 8: Comparing different replay buffer compositions in the class incremental case. Each matrix compares the mean final test accuracy of 10 networks trained with different replay buffer compositions achieved by modifying Goldilocks’ $quick$ and $slow$ hyperparameters. The ($quick$, $slow$) = ($0\%$, $0\%$) case, marked by a black box, corresponds to uniform sampling. (a) Replicating Fig. 4(b) with class incremental learning (b) Replicating Fig. 6 with class incremental learning. In both cases, the results mirror the task-incremental case, although with lower performances.
(a) CIFAR-100-5
(b) CIFAR-10-5
(c) TinyImagenet-10
Figure 9: Goldilocks vs. uniform sampling with various continual learning methods, in class incremental settings. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error.
(a) CIFAR-100-2
(b) CIFAR-10-2
(c) TinyImagenet-2
(d) CIFAR-100-2
(e) CIFAR-10-2
(f) TinyImagenet-2
Figure 10: Extendeing Fig.2 to CIFAR-100-2, Cifar-10-2 and a subset of TinyImagenet-2. (top) The first task’s binary epoch-wise classification matrices $M$ for the test data of each dataset. Rows denote examples, and columns denote epochs, indicating correct classifications by a majority of the networks. Rows are sorted by learning speed. Faster-learned examples from the first task are less likely to be forgotten at the end of the second task. (bottom) The mean learning speed of examples in the first task vs. the percentage of networks that remember them at the end of the second task. Networks forget more examples learned slowly across all 3 datasets.
(a) CIFAR-10-2
Figure 11: Extending Fig. 3(d) to CIFAR-10-2. We plot the mean learning speed of remembered examples by $10$ models trained with different buffer sizes. Models with bigger replay buffers remember slower-to-learn examples.
(a) ResNet-18
(b) Small ResNet
(c) Tiny ResNet
Figure 12: Comparing the extended epoch-wise classification matrix for different architectures, rows represent examples, and columns represent epochs, indicating correct classifications by a majority of 10 networks. In (a), we train ResNet-18. In (b), we train ResNet-18 with both width and depth reduced by half. In (c), we train ResNet-18 with both width and depth reduced by a factor of four. In all cases, rows are sorted by learning speed. Examples learned quickly in the first task are less likely to be forgotten after the second task.
(a) DER
(b) DER++
(c) GEM
(d) AGEM
(e) RPC
(f) X-DER
Figure 13: Comparing different replay buffer compositions for various continual learning methods, with a buffer of $1\%$. The experimental setup replicates the one done in Fig. 3.2. Like the experience replay case, focusing on examples learned midway through the learning process is most beneficial. Due to the smaller buffer, removing more easier examples is better.
Appendix B Correlations between learning speed and catastrophic forgetting in different datasets
In Section 2, we established a link between the learning speed of examples and catastrophic forgetting, primarily utilizing ResNet-18 networks trained on CIFAR-100-2. Here, we expand upon these findings by demonstrating similar phenomena across different datasets, architectures, and task numbers.
Other Datasets. We first explore other datasets using the experimental setup detailed in Fig. 2. We extend our analysis to CIFAR-10-2, CIFAR-100-2, and a subset of TinyImagenet-2 (comprising the initial 40 classes). The extended epoch-wise classification matrices for each task and the correlation between learning speed and the percentage of networks remembering each example are plotted in Fig. 10. Consistently across all three datasets, we observe a robust correlation between learning speed and the likelihood of example retention during continual training. Notably, faster-learned examples exhibit lower rates of catastrophic forgetting, as evidenced by both quantitative correlations and visual inspection of the extended epoch-wise classification matrices.
Further, in Fig. 11, we expand upon Fig.3 to include CIFAR-10-2. This extension reveals that as the buffer size increases, continual models tend to retain slower-to-learn examples. Analogous to the results observed in CIFAR-100-2, as depicted in Fig.3, we also observe a near-perfect negative correlation between the buffer size and the mean learning speed of remembered examples in CIFAR-10-2.
Other Architectures. We extend the results from Fig. 2 to architectures beyond ResNet-18 by creating two new, smaller variants: Small-ResNet and Tiny-ResNet. These are formed by reducing both the width and depth of ResNet-18 by a factor of 2 and 4 respectively. We replicate the experimental setup from Fig.2, training 10 models of each architecture on CIFAR-100-2 without a replay buffer, and storing the extended epoch-wise classification matrix. Each matrix row is sorted by the learning speed of the examples and plotted in Fig. 12. All matrices depict the classification of the test dataset. In all architectures, learning speed is highly correlated with catastrophic forgetting: networks forget more examples learned later in training while retaining an almost perfect recollection of those learned early on. Additionally, stronger architectures, similar to larger buffers, can remember slower-to-learn examples. This reinforces the motivation for Goldilocks, as it shows that the examples selected by Goldilocks are those the model could learn independently if it had a stronger architecture.
Other continual learning algorithms. In §3.4, we demonstrated the benefits of non-uniform buffer compositions for various continual learning methods. We extend these findings, initially shown for GEM in Fig. 6, to additional algorithms including DER, DER++, GEM, A-GEM, RPC, and X-DER. For each algorithm, we trained 10 networks with different buffer compositions by varying the $quick$ and $slow$ hyperparameters of Goldilocks. For each algorithm, we trained 10 networks on different buffer compositions, achieved by varying the $quick$ and $slow$ hyperparameters of Goldilocks. In all cases, we used the hyperparameters and architectures suggested in the original works. We used a buffer of size $500$ ($1\%$ of the data), due to its popularity in previous works. Consistent with the GEM results, we found that replay buffers focused on examples learned mid-way through the training process were most beneficial. Additionally, in all cases, removing slower-to-learn examples proved advantageous, given the small buffer size. These results can be found in Fig. 13.
Different number of tasks.
In most experiments presented in the paper, we focused on continual learning with 2 tasks. When considering a larger number of tasks, additional hyperparameters emerge, as the $quick$ and $slow$ parameters apply to each intermediate task. These can be tuned similarly to the two-task case, either through grid-search with a rotation task or using simple heuristics. To avoid excessive hyperparameter tuning when training on more than two tasks, we tuned the $quick$ and $slow$ values for the first task and kept them constant for all subsequent tasks. Results from training on CIFAR-10-5, CIFAR-100-5, and TinyImageNet-10 are shown in Fig. 17. We observe that Goldilocks improves over uniform sampling in these cases as well. Notably, further tuning of all Goldilocks hyperparameters could yield even better results.
Appendix C standard errors
In the main paper, we omitted the standard errors for visualization porpuses from Figs. 3.2,5,6 and Table 3.5. In both cases, these were usually very small and did not affect the qualitative results. For completeness, we report the omitted values in this section. The standard error for Figs. 3.2,5,6 can be found in Figs. C,15,16 respectively. A replication of Table 3.5, with the standard errors, can be found in Table C.
(a) CIFAR-100-2, $4\%$ buffer
(b) CIFAR-100-2, $40\%$ buffer
(c) TinyImagenet-2, $20\%$ buffer
Figure 14: Standard errors for Fig. 3.2. The error is taken over 10 repetitions in each experiment.
(d) (A)–>(B)
(e) (A)–>(C)
Figure 15: Standard errors for Fig. 5. The error is taken over 10 repetitions in each experiment.
(a) GEM, CIFAR-100-2, $1\%$ buffer
Figure 16: Standard errors for Fig. 6. The error is taken over 10 repetitions in each experiment.
Table 2: Same as Table 3.5, but with the addition of standard errors. Each entry is the mean test accuracy across all tasks of 10 networks trained using experience replay on CIFAR-100-2. Bold denotes the method with the highest accuracy.
CIFAR-10-2 CIFAR-100-2
Buffer size $4\%$ $40\%$ $4\%$ $40\%$
Random 87.67 $\pm$ 0.06 95.03 $\pm$ 0.05 69.03 $\pm$ 8.89 79.24 $\pm$ 1.82
Max entropy 84.44 $\pm$ 0.23 94.63 $\pm$ 0.06 64.27 $\pm$ 2.11 77.91 $\pm$ 0.91
IPM 85.9 $\pm$ 0.16 94.76 $\pm$ 0.13 64.16 $\pm$ 4.13 75.73 $\pm$ 1.43
GSS 85.28 $\pm$ 0.13 95.55 $\pm$ 0.08 64.15 $\pm$ 5.01 80.11 $\pm$ 0.81
Herding 88.41 $\pm$ 0.09 94.53 $\pm$ 0.12 71.07 $\pm$ 9.49 78.97 $\pm$ 0.1
Goldilocks 88.99 $\pm$ 0.03 96.03 $\pm$ 0.1 71.53 $\pm$ 7.31 80.62 $\pm$ 1.11
(a) CIFAR-100-5
(b) CIFAR-10-5
(c) TinyImagenet-10
Figure 17: Goldilocks vs. uniform sampling with various continual learning methods. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error. (a) CIFAR-100-5 with a buffer size of $5\%$. (b) CIFAR-10-5 with a buffer of size $5\%$ (c) TinyImagenet-10 with a buffer of size $10\%$. Goldilocks significantly improves performance across all methods and datasets.
(a) CIFAR-100 –> CIFAR-100
(b) CIFAR-100 –> rotated
(c) CIFAR-100 –> random labels
Figure 18: Comparing replay buffer compositions when training the same original task, followed by different subsequent tasks. Each entry in the matrices denotes the mean final test accuracy of $10$ networks, trained with experience replay on 2 tasks, with a buffer size of $40\%$. This Figure replicated the same experimental setup as Fig. 5. In (a) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100. In (b) we train the models on the 50 first classes of CIFAR-100, followed by a rotation classification task on the same examples (see text). In (c) we train the models on the 50 first classes of CIFAR-100, followed by the last 50 classes of CIFAR-100, but with random labels. In all cases, the same buffer compositions of the first task give the same qualitative results, suggesting that the optimal composition of the buffer of the first task is independent of the subsequent task.
Appendix D Independency of the optimal buffer composition when changing the subsequent tasks
We empirically examine how subsequent tasks impact the ideal composition of a replay buffer for a given task. We find that regardless of the similarity or dissimilarity between subsequent tasks and the original task, the optimal replay buffer composition remains largely independent and consistent. To show this, we conduct experiments with three distinct tasks denoted A, B, and C. By training the model on task A and subsequently introducing either task B or task C while keeping the original task unchanged, we analyze the ideal replay buffer composition under varying subsequent tasks.
In Section 3.3, we demonstrate that when selecting tasks A, B, and C from the same dataset, the optimal buffer composition of task A remains unaffected by tasks B or C. We observe consistent quantitative behavior across different $quick$ and $slow$ parameters. We further explore task variations by setting A as the first 50 classes of CIFAR-100, B as the last 50 classes of CIFAR-100, and C as a rotation classification [13] task. In the rotation classification task, examples from A are randomly rotated by angles of $\{90\degree,180\degree,270\degree\}$, with labels adjusted accordingly. We evaluate the performance of 10 networks trained continuously with different buffer compositions first on task A and then on task B (Fig.18(a)), and on task A followed by task C (rotation) (Fig.18(b)). Consistent with previous findings, the performance of various buffer compositions remains consistent, independent of the choice of B or C. This property enables us to utilize the rotation task for hyper-parameter grid search without requiring additional label data.
We further modify task C to comprise examples from the last 50 classes of CIFAR-100 but with random labels [57]. Training on random labels forces the network to memorize the examples, as no generalization is possible, eliminating any potential overlap between the subsequent tasks. The results of training on task A followed by the random label task are depicted in Fig. 18(c). Analogous to the rotation classification scenario, we observe that the same buffer compositions of task A remain effective, further suggesting that the optimal buffer composition of A does not depend on the subsequent task.
Appendix E Goldilocks with different continual learning methods
Here, we compare Goldilocks to uniform sampling with different continual learning algorithms. We tested several different algorithms on CIFAR-100-5, CIFAR-10-5, and TinyImagenet-10. The results, which extend Fig. 7, are plotted in Fig. 17.